# Smooth Tchebycheff Scalarization for Multi-Objective Optimization

### v2

@@ C.2.3 Hypervolume Definition
Appendix D Additional Experimental Studies
D.1 Office-31
@@
Full results on the Office-31 dataset are shown in Table 5. We run the experiments with MoCO, TCH, and STCH methods by ourselves, and the rest results are from Lin et al. (2023) with the same setting. Our proposed STCH scalarization method performs the best on the Amazon task, and achieves the second best performance on DSLR and Webcam, which leads to the best average performance and the best $\Delta_{p}$. In fact, it is the only method that can dominate the STL baseline on all tasks.
D.2 QM9
@@
$\mu$ $\alpha$ $\epsilon_{\text{HOMO}}$ $\epsilon_{\text{LUMO}}$ $\langle R^{2}\rangle$ ZPVE $U_{0}$ $U$ $H$ $G$ $c_{v}$ $\Delta_{p}\uparrow$
@@
Full results on the QM9 dataset are shown in Table 6. We run the experiments with TCH and STCH by ourselves, and the other results are from Lin et al. (2023). Our STCH scalarization has a promising second best overall performance, while the gap to the best DB-MTL (Lin et al., 2023) is very tight. It should be noticed that DB-MTL is an adaptive gradient method that requires a much longer runtime as reported in Table 4. Therefore, our proposed STCH scalarization method can serve as a very promising alternative for solving MTL problems.
D.3 Efficient Pareto Set Learning
Table 7: The runtime for learning the Pareto set of the four bar truss design problem with different methods.
@@ D.3 Efficient Pareto Set Learning
Runtime 6s (1x) 6s (1x) 268s (44.6x) 6s (1x) 6s (1x)
We report the runtimes of different methods to learn the Pareto set of the four bar truss design problem in Table 7. According to the results, our proposed STCH scalarization method share a similar runtime with other scalarization methods (e.g., LS, COSMOS, and TCH), while the gradient-based EPO method takes significantly longer runtime ($44.6$x) to learn the Pareto set. We also report the log hypervolume difference ($\log(\Delta HV)$) of different methods during the optimization process for $11$ different multi-objective optimization problems in Figure 6. It is clear that our proposed STCH scalarization method has the best overall performance on Pareto set learning.

### v3

@@ D.2.3 Hypervolume Definition
Appendix E Additional Experimental Studies
E.1 Office-31
@@
Full results on the Office-31 dataset are shown in Table 6. We run the experiments with MoCO, TCH, and STCH methods by ourselves, and the rest results are from Lin et al. (2023a) with the same setting. Our proposed STCH scalarization method performs the best on the Amazon task, and achieves the second best performance on DSLR and Webcam, which leads to the best average performance and the best $\Delta_{p}$. In fact, it is the only method that can dominate the STL baseline on all tasks.
E.2 QM9
@@
$\mu$ $\alpha$ $\epsilon_{HOMO}$ $\epsilon_{LUMO}$ $\langle R^{2}\rangle$ ZPVE $U_{0}$ $U$ $H$ $G$ $c_{v}$ $\Delta_{p}\uparrow$
Single-Task Baseline
@@
Full results on the QM9 dataset are shown in Table 7. We run the experiments with TCH and STCH by ourselves, and the other results are from Lin et al. (2023a). Our STCH scalarization has a promising second best overall performance, while the gap to the best DB-MTL (Lin et al., 2023a) is very tight. It should be noticed that DB-MTL is an adaptive gradient method that requires a much longer runtime as reported in Table 4. Therefore, our proposed STCH scalarization method can serve as a very promising alternative for solving MTL problems.
E.3 More Comparison to Linear Scalarization
Table 8: Results on the CelebA dataset.
Average Task Accuracy
Unit. Scal. 9.090e-01$\pm$7.568e-04
IMTL 9.093e-01$\pm$7.631e-04
MGDA 9.022e-01$\pm$9.687e-04
GradDrop 9.098e-01$\pm$3.383e-04
PCGrad 9.093e-01$\pm$1.108e-03
RLW Diri. 9.099e-01$\pm$7.845e-04
RLW Norm. 9.095e-01$\pm$1.012e-03
STCH (Ours) 9.098e-01$\pm$4.692e-04
We also test our proposed STCH scalarization on the CelebA dataset with the same setting as in Kurin et al. (2022). We independently run STCH $3$ times and report the mean of average task accuracy and standard deviation, while all other results are directly from Kurin et al. (2022). STCH has a simple equal weight for all tasks which is not tuned. According to the results in Table 8, STCH achieves similar performance with most other methods, and significantly outperforms MGDA. We find that STCH achieves almost zero training loss for all tasks, which is consistent with most other methods as reported in Kurin et al. (2022). In this case, the testing accuracy mainly depends on the generalization performance rather than the optimization performance (on training loss). Since the STCH scalarization is developed from the viewpoint of multi-objective optimization, we currently do not have a straightforward way to analyze its generalization performance. Recently, Chen et al. (2023) have investigated the generalization gap for the stochastic adaptive gradient method. We leave the generalization analysis for STCH in the stochastic setting to future work.
On the other hand, our proposed STCH scalarization is lightweight, with a computational overhead similar to that of linear scalarization. It is also interesting to further investigate its performance with tuned weights and/or regularization for multi-task learning as in Kurin et al. (2022) and Xin et al. (2022).
E.4 Multi-Objective Bayesian Optimization
Table 9: Results (hypervolume difference $\Delta\text{HV}\downarrow$) for multi-objective Bayesian optimization on $6$ synthetic benchmark problems and $5$ real-world engineering design problems.
F1 F2 F3 F4 F5 F6 BarTruss PressureVessel DiskBrake GearTrain RocketInjector
TCH 0.0183 0.0185 0.0413 0.0891 0.0602 0.0917 0.0429 0.0441 0.0262 0.0262 0.0449
STCH 0.0147 0.0110 0.0304 0.0476 0.0455 0.0575 0.0233 0.0194 0.0214 0.0198 0.0332
In this subsection, we test our proposed STCH on multi-objective Bayesian optimization. We follow the setting proposed in Lin et al. (2022c), where a Pareto set learning approach with TCH scalarization (PSL-TCH) has been proposed for multiobjective Bayesian optimization. We simply replace the TCH scalarization in PSL-TCH with the proposed STCH scalarization to obtain a new method PSL-STCH. We then compare the new PSL-STCH with PSL-TCH on different synthetic and multiobjective engendering design problems with a small evaluation budget (i.e. 100 evaluations) as in Lin et al. (2022c). According to Table 9 with the hypervolume difference ($\Delta HV$) metrics, PSL-STCH can outperform PSL-TCH on all problems. This result confirms that the proposed STCH scalarization can serve as an off-the-shelf approach for different multiobjective optimization applications, including Bayesian optimization.
E.5 Efficient Pareto Set Learning
[…]
LS 0.8602 0.8629 0.7826 0.3172 0.3712 0.2892 0.8805 1.1633 0.9738 1.0169 0.7051
COSMOS 0.8608 0.8614 0.8638 0.5283 0.5300 0.5242 0.8803 1.1425 0.9710 1.0174 0.8091
EPO 0.8653 0.8689 0.8564 0.5324 0.5349 0.5236 0.8772 1.1592 0.9805 1.0175 0.7889
TCH 0.8676 0.8686 0.8582 0.5345 0.5364 0.5287 0.8795 1.1611 0.9765 1.0170 0.8198
STCH 0.8707 0.8709 0.8670 0.5365 0.5372 0.5316 0.8829 1.1632 0.9864 1.0177 0.8363
Hypervolume Metrics
For the efficient Pareto set learning problem, we report the hypervolume for all problems in Table 10. For each problem, we first normalize the objective value of all solutions for all methods into $[0,1]$ with the same nadir point and ideal point, then calculate the hypervolume with reference point $[1.1,1.1]^{m}$ where $m$ is the number of objectives. Our proposed STCH achieves the best overall performance on these problems.
Table 11: The run time for learning the Pareto set of $5$ real-world enginnering design problems with different methods.
@@ E.5 Efficient Pareto Set Learning
Bar Truss 6s (1x) 6s (1x) 268s (45x) 6s (1x) 6s (1x)
Hatch Cover 6s (1x) 6s (1x) 277s (46x) 6s (1x) 6s (1x)
Disk Brake 9s (1x) 9s (1x) 329s (37x) 9s (1x) 9s (1x)
Gear Train 9s (1x) 9s (1x) 337s (38x) 9s (1x) 9s (1x)
Rocket Injector 9s (1x) 9s (1x) 324s (36x) 9s (1x) 9s (1x)
Run time
In addition, we also report the runtimes of different methods to learn the Pareto set for five multi-objective engineering design problems in Table 11. According to the results, our proposed STCH scalarization method share a similar runtime with other scalarization methods (e.g., LS, COSMOS, and TCH), while the gradient-based EPO method takes significantly longer runtime ($36$x to $44$x) to learn the Pareto set.